Solve the systems of equations by using the substitution method. (If the system is dependent, enter a general solution in terms of c. If there is no solution, enter NO SOLUTION.)
{5x− 4y = −17
y= 5x − 7
(x, y) =

Answer by math_helper(2461)


5x-4y = -17
y = 5x-7

Substituting for y in the top equation by the RHS of the bottom equation:
5x - 4(5x-7) = -17
5x - 20x + 28 = -17
-15x = -45
15x = 45
x = 45/15 = 3 —> y = 5(3)-7 = 15-7 = 8

Ans: x=3, y=8